# Classical and quantum cyclic redundancy check codes

### v2

@@ Abstract
We prove that certain classical cyclic redundancy check codes can be used for classical error correction and not just classical error detection. We extend the idea of classical cyclic redundancy check codes to quantum cyclic redundancy check codes. This allows us to construct quantum stabiliser codes which can correct burst errors where the burst length attains the quantum Reiger bound. We then consider a certain family of quantum cyclic redundancy check codes for which we present a fast linear time decoding algorithm.
@@ 1 Introduction
In many classical channels, including communication channels and storage mediums, errors tend to come in bursts, i.e. located in a short interval. There are classical codes specifically designed to deal with these types of errors. Cyclic redundancy check (CRC) codes are a class of commonly employed codes which are used to detect burst errors. These codes are cyclic codes and have the ability to detect cyclic burst errors of maximum length set out by the Reiger bound. In this article we will also show that, if chosen carefully, there are CRC codes which can also correct burst errors of maximum length set out by the Reiger bound. Analogously, in the quantum world, quantum errors can also be correlated in space and time [4] and quantum channels can introduce errors which are localised, i.e. quantum burst errors. Another objective of this article is to construct quantum CRC codes. We show that there are quantum CRC codes which meet the quantum Reiger bound. The quantum Reiger bound was proven in [5, Theorem 3], where some constructions of codes meeting the bound were given.
@@
We say that $C$ can correct burst errors of length $t$ if we can identify $e$, given any $u+e$, where $u\in C$ and $e$ is any vector of ${\mathbb{F}}_{q}^{n}$ whose non-zero coordinates lie in a burst $\{i,\ldots,i+t-1\}$, where the coordinate positions are read modulo $n$. If $t$ is minimal with this property then we say that $e$ has burst weight $t$. Note that we could restrict to non-cyclic bursts where the coordinates are not read modulo $n$ but this would be more restrictive, so the codes we will construct will always be able to correct cyclic bursts as well.
@@ 2 Quantum Error Correction
for all $\bra{\psi_{i}}\ket{\psi_{j}}=\delta_{ij}$ and for all $E,E^{\prime}\in\mathcal{E}$, where |$\ket{\psi_{i}}$ and |$\ket{\psi_{j}}\in\mathcal{Q}$, and $a(E,E^{\prime})$ is a constant which depends only on $E$ and $E^{\prime}$, see [13, Theorem 10.1] or [2].
If condition (1) holds then we can correct any linear combination of the operators in $\mathcal{E}$, see [13, Theorem 10.2]. This observation is due to Gottesman [7].
We will restrict our quantum error correcting code to stabiliser codes over finite fields of prime order. Thus $\mathcal{Q}$ will be a subspace of $({\mathbb{C}}^{p})^{\otimes n}$, where $p$ is a prime.
@@
3 Stabiliser Codes and Additive Codes
@@
$((a|b),(a^{\prime}|b^{\prime}))_{s}=(a\cdot b^{\prime}-a^{\prime}\cdot b).$ (2)
@@
$MM^{\prime}=w^{(\tau(M),\tau(M^{\prime}))_{s}}M^{\prime}M$ (3)
@@
The centraliser of $S$ in $\mathcal{P}_{n}$ is
$\mathrm{Centraliser}(S)=\{E\in\mathcal{P}_{n}\ |\ EME^{-1}=M,\ \mathrm{for}\ \mathrm{all}\ M\in S\}.$
Since $S$ is abelian, the centraliser clearly contains
@@
is the image of the centraliser of $S$ under the map $\tau$.
@@
The stabiliser code associated with $S$ is
@@
All quantum error-correcting codes considered in this article will be stabiliser codes.
The following lemma is [13, Proposition 10.5].
@@
Condition (1) implies the following theorem for stabiliser codes, [13, Theorem 10.8].
@@
If $Q(S)$ is unable to correct $E_{1}^{\dagger}E_{2}$ then $E_{1}^{\dagger}E_{2}\in\mathrm{Centraliser}(S)\setminus[S]$. Therefore, $\tau(E_{1}^{\dagger}E_{2})=\tau(E_{1})+\tau(E_{2})=v_{1}+v_{2}\in C^{\perp_{a}}\setminus C$.
Thus, we have the following statement. If for any vectors $v_{1},v_{2}\in{\mathbb{F}}_{p}^{2n}$ of symplectic weight (respectively burst symplectic weight) at most $t$, the vector $v_{1}+v_{2}\not\in C^{\perp_{a}}\setminus C$ then $Q(S)$ is a quantum $t$-error correcting code (respectively quantum $t$-burst error correcting code).
@@
We will make use of the following particular construction of stabiliser codes. For any matrix $H$, we denote by $H_{\ell}$ the matrix $H$ in which all columns have been shifted cyclically $\ell$ positions to the right. Thus, we have $H_{0}=H$ and $H_{-\ell}$ is the matrix obtained from $H$ in which all columns have been shifted cyclically $\ell$ positions to the left.
@@
is the generator matrix of a linear code $C=\tau(S)$, for some abelian subgroup $S$ of $\mathcal{P}_{n}$ of size $p^{n-k}$ defining an $[\![n,k]\!]$ stabiliser code $\mathcal{Q}(S)$.
@@
then $s_{i}=(\tau(E),\tau(M_{i}))_{s}$, so the syndrome of $E$ is given by
@@ 4 Cyclic redundancy check (CRC) codes
The syndrome $s(e)=eH^{t}$ of CRC($g$) is $e(X)\ mod\ g$, where we identify the row vector $eH^{t}$ with a polynomial of degree (at most) $n-k$ in the obvious way.
@@
We say that $g$ has the c-property if furthermore no multiple of $g$ is the sum of two polynomials of burst weight at most $\lfloor\frac{n-k}{2}\rfloor$. Observe that this is equivalent to choosing $g$ so that the $k$-dimensional cyclic code generated by $g$ contains no codewords which are the sum of two vectors of burst weight at most $\lfloor\frac{n-k}{2}\rfloor$.
@@
If $s(e_{1})=s(e_{2})$ then $s(e_{1}-e_{2})=0$. By Lemma 8, $e_{1}-e_{2}$ is a multiple of $g$. By assumption, both $e_{1}$ and $e_{2}$ have burst weight at most $\lfloor\frac{n-k}{2}\rfloor$ which contradicts the c-property.
@@
Tables 4, 5, 6, 7 list all polynomials $g\in{\mathbb{F}}_{2}[X]$ which are factors of $X^{n}-1$ and have the c-property for all $n\leqslant 39$. We have omitted $g(X)=(X^{n}-1)/(X-1)$ and $g(X)=X-1$ from the table since these always have the c-property.
@@ 5 Quantum CRC codes
by identifying $e_{1},e_{2}\in{\mathbb{F}}_{2}^{n}$ with a polynomial of degree at most $n-1$.
@@
Suppose that $e(X)=e_{1}(X)+X^{n}e_{2}(X)$ is a polynomial of burst weight at most $2\ell$.
We will prove that $s_{Q}(e)=0$ then $e=0$.
@@
$X^{-l}f_{1}+X^{l}f_{1}+f_{2}=0\pmod{X^{n}-1}$
or the burst length of $X^{-l}f_{1}+X^{l}f_{1}+f_{2}$ is at least $n-k+1$. But the burst length of $X^{-l}f_{1}+X^{l}f_{1}+f_{2}$ is at most $2\ell+\deg(f_{1})+1\leqslant 4\ell$. Hence,
$X^{-l}f_{1}+X^{l}f_{1}+f_{2}=0\pmod{X^{n}-1}.$
@@
Suppose that $e(X)=e_{1}(X)+X^{n}e_{2}(X)$ and $e^{\prime}(X)=e^{\prime}_{1}(X)+X^{n}e^{\prime}_{2}(X)$ are polynomials of burst weight at most $\ell$.
@@ 6 A fast decoding algorithm for a particular family of quantum CRC codes
If $g(x)=x^{n-k}+x^{n-2k}+\dots x^{k}+1$ then there is a decoding algorithm for $Q_{g}(S)$ whose complexity scales linearly with n.
@@
For example, if $m=9$ and $c=2$ then we have the $[\![9,1]\!]$ code given by the following matrix.
$\mathrm{G}_{[\![9,1]\!]}=\begin{pmatrix}\begin{array}[]{ccccccccc|ccccccccc}1&0&0&0&0&0&0&0&1&0&1&1&0&0&0&1&1&0\\ 0&1&0&0&0&0&0&0&1&0&1&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&0&0&1&1&1&0&0&1&0&1&0&0\\ 0&0&0&1&0&0&0&0&1&0&0&0&0&0&1&1&0&0\\ 0&0&0&0&1&0&0&0&1&0&1&1&0&0&0&0&0&0\\ 0&0&0&0&0&1&0&0&1&0&1&0&1&0&0&1&1&0\\ 0&0&0&0&0&0&1&0&1&0&1&0&0&1&0&1&0&1\\ 0&0&0&0&0&0&0&1&1&1&1&0&0&0&1&1&0&0\par\par\end{array}\end{pmatrix}$ (8)
If we look at this matrix we can clearly see 3 columns whose weight is more than $2$. The last column of the left-hand side matrix is the all $1$ vector, and the $c$ and $m-c$ column of the right-hand side are the all $1$ vector summed with a vector of weight one.
@@
A set of generators for the stabiliser group $S$ is given by the row of the matrix
@@
A set of generators for the stabiliser group given by this matrix is given by the rows of the matrix
@@ 8 Conclusion
In conclusion, in this article we have introduced a method to construct quantum CRC codes from classical CRC codes. We note that this method can also be used with any classical code to construct a quantum stabiliser code. The classical CRC codes which obtain the Reiger bound, will give quantum CRC codes which attain the quantum Reiger bound. Since there are classical CRC codes which obtain the Reiger bound for all length $n$ and dimension $k$, we conclude that the quantum Reiger bound is always attainable.

### v3

@@ Abstract
We prove that certain classical cyclic redundancy check codes can be used for classical error correction and not just classical error detection. We extend the idea of classical cyclic redundancy check codes to quantum cyclic redundancy check codes. This allows us to construct quantum stabilizer codes which can correct burst errors where the burst length attains the quantum Reiger bound. We then consider a certain family of quantum cyclic redundancy check codes for which we present a fast linear time decoding algorithm.
@@ 1 Introduction
In many classical channels, including communication channels and storage mediums, errors tend to come in bursts, i.e. located in a short interval. There are classical codes specifically designed to deal with these types of errors. Cyclic redundancy check (CRC) codes are a class of commonly employed codes which are used to detect burst errors. These codes are cyclic codes and have the ability to detect cyclic burst errors of maximum length set out by the Reiger bound. In this article we will also show that, if chosen carefully, there are CRC codes which can also correct burst errors of maximum length set out by the Reiger bound. Analogously, in the quantum world, quantum errors can also be correlated in space and time [5] and quantum channels can introduce errors which are localised, i.e. quantum burst errors. Another objective of this article is to construct quantum CRC codes. We show that there are quantum CRC codes which meet the quantum Reiger bound. The quantum Reiger bound was proven in [6, Theorem 3], where some constructions of codes meeting the bound were given.
@@
We say that $C$ can correct burst errors of length $t$ if we can identify $e$, given any $u+e$, where $u\in C$ and $e$ is any vector of ${\mathbb{F}}_{q}^{n}$ whose non-zero coordinates lie in a burst $\{i,\ldots,i+t-1\}$, where the coordinate positions are read modulo $n$. If $t$ is minimal with this property then we say that $e$ has burst weight $t$. Note that we could restrict to non-cyclic bursts where the coordinates are not read modulo $n$ but this would be more restrictive, so the codes we will construct will always be able to correct cyclic bursts as well. We also choose the modulo in the set $\{1,\ldots,n\}$ instead of $\{0,\ldots,n-1\}$, since we are referring to coordinates.
@@ 2 Quantum Error Correction
for all $\bra{\psi_{i}}\ket{\psi_{j}}=\delta_{ij}$ and for all $E,E^{\prime}\in\mathcal{E}$, where |$\ket{\psi_{i}}$ and |$\ket{\psi_{j}}\in\mathcal{Q}$, and $a(E,E^{\prime})$ is a constant which depends only on $E$ and $E^{\prime}$, see [14, Theorem 10.1] or [3].
If condition (1) holds then we can correct any linear combination of the operators in $\mathcal{E}$, see [14, Theorem 10.2]. This observation is due to Gottesman [8].
We will restrict our quantum error correcting code to stabilizer codes over finite fields of prime order. Thus $\mathcal{Q}$ will be a subspace of $({\mathbb{C}}^{p})^{\otimes n}$, where $p$ is a prime.
@@
3 Stabilizer Codes and Additive Codes
@@
$((a|b),(a^{\prime}|b^{\prime}))_{s}=a\cdot b^{\prime}-a^{\prime}\cdot b.$ (2)
@@
$MM^{\prime}=w^{(\tau(M),\tau(M^{\prime}))_{s}}M^{\prime}M.$ (3)
@@
The centralizer of $S$ in $\mathcal{P}_{n}$ is
$\mathrm{Centralizer}(S)=\{E\in\mathcal{P}_{n}\ |\ EME^{-1}=M,\ \mathrm{for}\ \mathrm{all}\ M\in S\}.$
Since $S$ is abelian, the centralizer clearly contains
@@
is the image of the centralizer of $S$ under the map $\tau$.
@@
The stabilizer code associated with $S$ is
@@
All quantum error-correcting codes considered in this article will be stabilizer codes.
The following lemma is [14, Proposition 10.5].
@@
Condition (1) implies the following theorem for stabilizer codes, [14, Theorem 10.8].
@@
If $Q(S)$ is unable to correct $E_{1}^{\dagger}E_{2}$ then $E_{1}^{\dagger}E_{2}\in\mathrm{Centraliser}(S)\setminus[S]$. Therefore, $\tau(E_{1}^{\dagger}E_{2})=\tau(E_{1})+\tau(E_{2})=v_{1}+v_{2}\in C^{\perp_{s}}\setminus C$.
Thus, we have the following statement. If for any vectors $v_{1},v_{2}\in{\mathbb{F}}_{p}^{2n}$ of symplectic weight (respectively burst symplectic weight) at most $t$, the vector $v_{1}+v_{2}\not\in C^{\perp_{s}}\setminus C$ then $Q(S)$ is a quantum $t$-error correcting code (respectively quantum $t$-burst error correcting code).
@@
We will make use of the following particular construction of stabilizer codes. For any matrix $H$, we denote by $H_{\ell}$ the matrix $H$ in which all columns have been shifted cyclically $\ell$ positions to the right. Thus, we have $H_{0}=H$ and $H_{-\ell}$ is the matrix obtained from $H$ in which all columns have been shifted cyclically $\ell$ positions to the left.
@@
is the generator matrix of a linear code $C=\tau(S)$, for some abelian subgroup $S$ of $\mathcal{P}_{n}$ of size $p^{n-k}$ defining an $[\![n,k]\!]$ stabilizer code $\mathcal{Q}(S)$.
@@ 4 Syndrome of a stabilizer code
then $s_{i}=(\tau(M_{i}),\tau(E))_{s}$, so the syndrome of $E$ is given by
@@ 5 Cyclic redundancy check (CRC) codes
For any vector $v=(v_{1},\ldots,v_{n})$, we associate a polynomial which we write as
$v(X)=\sum_{i=0}^{n}v_{i}X^{i-1}$
and vice-versa.
@@
The syndrome $s(e)=eH^{t}$ of CRC($g$) is $v$ where $v(X)=e(X)\ mod\ g$.
@@
We now introduce a new property for polynomials called the c-property. We say that $g$ has the c-property if no multiple of $g$ in $\mathbb{F}_{p}[X]/(X^{n}-1)$ is the sum of two polynomials of (cyclic) burst weight at most $\lfloor\frac{n-k}{2}\rfloor$. Observe that this is equivalent to choosing $g$ so that the $k$-dimensional cyclic code generated by $g$ contains no codewords which are the sum of two vectors of burst weight at most $\lfloor\frac{n-k}{2}\rfloor$.
@@
Let $e_{1}$ and $e_{2}$ be errors of burst weight at most $\lfloor\frac{n-k}{2}\rfloor$. If $s(e_{1})=s(e_{2})$ then $s(e_{1}-e_{2})=0$. By Lemma 8, $(e_{1}-e_{2})(X)$ is a multiple of $g$, which contradicts the c-property. Thus, $s(e_{1})\neq s(e_{2})$.
@@
Tables 4, 5,list all polynomials $g\in{\mathbb{F}}_{2}[X]$ which are factors of $X^{n}-1$ and have the c-property for all $n\leqslant 27$. We have omitted $g(X)=(X^{n}-1)/(X-1)$ and $g(X)=X-1$ from the table since these always have the c-property.
In [1], some analysis was carried out showing that decoding algorithms such as GRAND, can be used to correct errors using CRC codes. These results do not make use of polynomials which necessarily have the c-property. It would be interesting to have more sophisticated decoding algorithms which were tailored to a particular choice of $g$ and make use of Theorem 11.
@@ 6 Quantum CRC codes
by identifying $e_{1},e_{2}\in{\mathbb{F}}_{2}^{n}$ with a polynomial of degree at most $n-1$ as before.
@@
Suppose that $e(X)=e_{1}(X)+X^{n}e_{2}(X)$ is a polynomial of symplectic burst weight at most $2\ell$.
We will prove that if $s_{Q}(e)=0$ then $e=0$.
@@
$X^{-\ell}f_{1}+X^{\ell}f_{1}+f_{2}=0\pmod{X^{n}-1}$
or the burst length of $X^{-\ell}f_{1}+X^{\ell}f_{1}+f_{2}$ is at least $n-k+1$. But the burst length of $X^{-\ell}f_{1}+X^{\ell}f_{1}+f_{2}$ is at most $2\ell+\deg(f_{1})+1\leqslant 4\ell$. Hence,
$X^{-\ell}f_{1}+X^{\ell}f_{1}+f_{2}=0\pmod{X^{n}-1}.$
@@
Suppose that $e(X)=e_{1}(X)+X^{n}e_{2}(X)$ and $e^{\prime}(X)=e^{\prime}_{1}(X)+X^{n}e^{\prime}_{2}(X)$ are polynomials of symplectic burst weight at most $\ell$.
@@ 7 A fast decoding algorithm for a particular family of quantum CRC codes
If $g(X)=X^{n-k}+X^{n-2k}+\dots+X^{k}+1$ then there is a decoding algorithm for $Q_{g}(S)$ whose complexity scales linearly with n.
@@
If we look at this matrix, see for example (10), we can clearly see 3 columns whose weight is more than $2$. The last column of the left-hand side matrix is the all $1$ vector, and the $c$ and $m-c$ column of the right-hand side are the all $1$ vector summed with a vector of weight one.
@@
A set of generators for the stabilizer group $S$ is given by the row of the matrix
@@
A set of generators for the stabilizer group given by this matrix is given by the rows of the matrix
@@ 9 Conclusion
In conclusion, in this article we have introduced a method to construct quantum CRC codes from classical CRC codes. We note that this method can also be used with any classical code to construct a quantum stabilizer code. The classical CRC codes which obtain the Reiger bound, will give quantum CRC codes which attain the quantum Reiger bound. Since there are classical CRC codes which obtain the Reiger bound for all length $n$ and dimension $k$, we conclude that the quantum Reiger bound is always attainable.